Find the measure of angle B if seven times the complement of angle B exceeds twice the supplement of angle B by 220.
edna_albaracin@yahoo.com.ph

Answer by mananth(16949)

let the measure of angle B be x deg
the complement will be 90-x
7 time scomplement = 7*(90-x)
..
The supplement will be 180-x
twice the supplement =2*(180-x)
..
7*(90-x)-2*(180-x)=220
630-7x-360+2x=220
-5x= 220+360-630
-5x=-50
x= -50/-5
x=10 deg